### v1

# Social Choice for AI Alignment: Dealing with Diverse Human Feedback

Abstract
Foundation models such as GPT-4 are fine-tuned to avoid unsafe or otherwise problematic behavior, so that, for example, they refuse to comply with requests for help with committing crimes or with producing racist text. One approach to fine-tuning, called reinforcement learning from human feedback, learns from humans’ expressed preferences over multiple outputs. Another approach is constitutional AI, in which the input from humans is a list of high-level principles. But how do we deal with potentially diverging input from humans? How can we aggregate the input into consistent data about “collective” preferences or otherwise use it to make collective choices about model behavior? In this paper, we argue that the field of social choice is well positioned to address these questions, and we discuss ways forward for this agenda, drawing on discussions in a recent workshop on Social Choice for AI Ethics and Safety held in Berkeley, CA, USA in December 2023.
Machine Learning, ICML
1 Introduction
Over the past year, reinforcement learning from human feedback (RLHF) has played a key role in making large language models (LLMs) more capable and controllable (Christiano et al., 2017; Ziegler et al., 2019). RLHF is now the primary strategy that leading AI companies such as OpenAI (OpenAI, 2023), Anthropic (Anthropic, 2023), Meta (Meta, 2023), and Google (Google, 2023) use to align pretrained LLM models with human values. However, RLHF faces many limitations and concrete challenges (Casper et al., 2023; Lambert & Calandra, 2023), including unrepresentative data (Prabhakaran et al., 2021; Feffer et al., 2023), unrealistic models of human decision-making (Hong et al., 2022; Freedman et al., 2021; Siththaranjan et al., 2023; Lambert et al., 2023), and insufficient modeling of human diversity (Kirk et al., 2023; Freedman et al., 2023). We hold the position that core ideas from social choice theory (Arrow, 2012; Fishburn, 1973; Kelly, 1988; Brandt et al., 2015)—primarily concerning whose preferences should be integrated into decisions and how this should be done—are needed to solve many of the open problems facing RLHF.
While models that are solely pretrained on internet data may produce repetitive or harmful text, RLHF enables training models to follow instructions (Ouyang et al., 2022) and produce helpful and “harmless” outputs (Bai et al., 2022a) based on human judgments. RLHF gathers example outputs from an LLM that has been pretrained to predict a text corpus. Next, humans are asked to select the outputs that best meet specified criteria (such as being “helpful” or “unbiased”). Humans may also manually write the outputs to be compared, but due to cost, human input is often limited to these comparative judgements. These judgments, often called preferences, are then used to fine-tune the LLM to produce more desirable outputs. From a social choice perspective, this method raises several critical questions: Which humans are asked to judge models? What criteria do they use? How are their judgments combined? And how do their expressed judgments relate to their actual preferences?
Constitutional AI (CAI), which involves reinforcement learning from AI feedback (RLAIF), is an alternate approach that directly addresses some of these questions (Bai et al., 2022b). Humans produce a “constitution” that explicitly specifies principles to guide the LLM training process. The LLM is then trained to align with this constitution. However, we must still decide who has input on the constitution and how it is constructed. Bai et al. (2022b) construct their constitution “in a fairly ad-hoc way […] for research purposes”, but developing safe and ethical AI requires a more principled approach, as exemplified in Ganguli et al. (2023) or announced in OpenAI (2024). How then should one aggregate diverse preferences into a representative constitution?
Social choice theory has long studied similar questions, and by taking into account its lessons, one can avoid making naïve mistakes and reinventing the wheel. In this paper, we argue that tools and theories from social choice should be applied to these open problems, in particular in RLHF, to help bridge challenging design problems to sociotechnical questions (Dobbe et al., 2021). Specifically, we demonstrate how such tools can be used to begin addressing which humans should provide input or feedback, what type of feedback they should provide, and how that feedback should be aggregated and used. We also highlight areas in which new work is required to extend social choice to new problems unique to training safe and ethical AI.
There are a number of advantages to addressing these problems in a principled way. First, it is likely to result in a fairer system that takes into account the input or feedback of a broader group of people. Second, there are reasons to believe that this will result in generally more accurate feedback about questions of truthfulness. Indeed, there is a significant body of literature on “epistemic democracy”—voting to settle questions about facts (Pivato, 2017). Intuitively, having input from a more diverse group of people makes it less likely that something important is missed. Third, it will likely result in broader buy-in into the system. For example, important issues such as political biases of LLMs (Motoki et al., 2023) have been hypothesized to emerge from the finetuning phase that follows pretraining (Rozado, 2024).
One may also have concerns about this approach; for example, is feedback from a diverse group of people going to be inconsistent and consequently result in inconsistent behavior from the system? Social choice theory provides a number of examples where naïve aggregation of preferences or judgments leads to choices in the aggregate that seem irrational, such as cyclical preferences (Schwartz, 2018) or logically inconsistent conclusions (List & Pettit, 2002). Then again, social choice theory also provides the tools for thinking about such issues and preventing them.
The field of social choice is not new to computer scientists; computational social choice (Brandt et al., 2015) is by now a well-studied topic, with a dedicated biennial workshop since 2006. However, while many of the researchers in this area affiliate with the AI community, there has not yet been much work connecting computational social choice to the alignment of modern AI systems.
In the remainder of this paper, we first give background on value alignment, RLHF, and social choice. Then we discuss a number of questions at the intersection of these topics. We believe that significant further research is required to answer each of these questions well and that good answers to them are needed to build AI systems in a responsible way based on potentially diverging feedback from multiple stakeholders. In contrast, ad-hoc approaches to these questions may result in systems that fail to represent their stakeholders well, that fail to address important issues, that marginalize significant groups of stakeholders, and that create a basis for conflict between groups of people or the multiple AI systems that represent them.
2 Background
Our proposed research agenda requires background on topics that have so far been studied by mostly disjoint communities. A reader familiar with some of these topics can skip the corresponding subsections.
2.1 Value Alignment
As advanced AI systems become increasingly capable, it becomes critical that they act in a way that aligns with human and societal values (Gabriel, 2020). There are many approaches to value alignment, including theoretical work to define formal games that AI agents must align with humans to solve (Shah et al., 2020), empirical investigation of the relationship between neural network activations and morally relevant output features (Zou et al., 2023), and evaluations of the ethical behavior of state-of-the-art models (Pan et al., 2023). RLHF is a particularly popular approach to value alignment, but it faces many limitations in its current form.
2.2 Reinforcement Learning from Human Feedback
Preference data collection
The first step in RLHF is to generate and evaluate a dataset of model outputs $\mathcal{Y}$. In vanilla RLHF, humans are then shown paired completions $\{y_{0},y_{1}\}\in\mathcal{Y}\times\mathcal{Y}$ to prompts $x\in\mathcal{X}$ of these outputs and asked to select which output $p\in\{y_{0},y_{1}\}$ they prefer from each pair (Christiano et al., 2017; Lee et al., 2021). Other RLHF variants require humans to rank or provide scores for groups of outputs (Ziegler et al., 2019; Ouyang et al., 2022), and many additional variations exist (Wu et al., 2023).
Reward model training
The next step is to fit a parameterized reward model $\varrho_{\theta}\>:\>\mathcal{Y}\rightarrow\mathbb{R}$. For LLMs, the reward model is typically a neural network with weights $\theta$. RLHF methods assume that there is a ground-truth reward function $\varrho_{\theta^{*}}$ that the human preferences reflect up to probabilistic noise. The reward model is then optimized to match the likelihoods of the human preferences observed in the data. If the training data comes from diverse sources, this implicitly amounts to a rather intransparent form of preference aggregation (Siththaranjan et al., 2023).
Optimizing the policy with RL
The final step is to use reinforcement learning to train a policy that maximizes rewards from the reward model. This involves many design decisions—which RL algorithm to use, how to regularize the updates, and whether to gather further online feedback during training. See Uc-Cetina et al. (2023) for a survey of methods and limitations for using RL to train LLMs.
2.3 Alternate Preference-based Fine-tuning Objectives
Due to the instability of popular RL optimizers such as Proximal Policy Optimization (PPO) (Schulman et al., 2017), several novel techniques for optimizing a model based on a collected preference dataset have been proposed. Rafailov et al. (2023) introduce Direct Preference Optimization (DPO), which recasts RLHF to converge to the optimum reward modeling solution by directly optimizing a loss on the preference label dataset, rather than sampling online from the LLM policy or training an explicit reward model. Another variant emerged to remove the dependency on pairwise data. Ethayarajh et al. (2023) propose a loss function termed Kahneman-Tversky Optimization (KTO) that enables learning a policy from unpaired preferences. The authors further claim that the effectiveness of various losses for RLHF depends on the properties they share with proposed human utility functions (Kahneman & Tversky, 1979).
2.4 Constitutional AI
Bai et al. (2022b) further explore the design space by introducing Constitutional AI (CAI), which relies on RL from AI Feedback (RLAIF). RLAIF is a set of techniques for using an AI model to augment or generate feedback data in the form of pairwise preferences or other signals (Lee et al., 2023; Sharma et al., 2024; Castricato et al., 2024). By employing a human-written set of principles, which they term a constitution, they use a separate LLM to generate artificial preference and instruction data that can be used for model fine-tuning. A constitution $\mathcal{C}$ is made up of a set of written principles $c_{i}$ that indicate specific aspects to focus on during a critique phase. The instruction data, which is largely out of the scope of this paper, is curated by repeatedly sampling a principle $c_{i}$ and asking the model to revise the current completion $y_{k}^{0}$ to the prompt $x_{k}$. This yields a series of instruction variants $\{y_{k}^{0},y_{k}^{1},\cdots,y_{k}^{n}\}$ from the principles $\{c_{i_{0}}^{0},c_{i_{1}}^{1},\cdots,c_{i_{n-1}}^{n-1}\}$ used for critique at each step. The final data point is the prompt $x_{k}$ with the final completion $y_{k}^{n}$, for some suitable $n$.
The preference data is constructed in a similar, yet simpler way by using a subset of principles from the constitution $\mathcal{C}$ as context for a feedback model. The feedback model is presented with a prompt $x$, a set of principles $\{c_{0},\cdots,c_{n}\}$, and two completions $y_{0}$ and $y_{1}$ labeled as answers (A) and (B) from a previous RLHF dataset. The feedback models’ probability of outputting either (A) or (B) is recorded as a training sample for the reward model, as discussed in Section 2.2.
2.5 Social Choice
Modern social choice theory began in the 1950s with Arrow’s Impossibility Theorem (Arrow, 1951). [^1] For its long prehistory, see McLean & Urken (1995). Arrow considered the problem of aggregating multiple individuals’ preferences—in the form of complete and transitive rankings of some set of alternatives—into a social preference, subject to a list of normative desiderata. In particular, Arrow assumed that the aggregation function should be defined for any family of individual preferences to be aggregated (Universal Domain); that the outputted social preference relation should be complete and transitive, like individual preferences, in which case the aggregation function is called a social welfare function; that the social preference between two alternatives $A$ and $B$ should depend only on individual preferences between $A$ and $B$ (Independence of Irrelevant Alternatives); and that unanimous individual preference for $A$ over $B$ should imply social preference for $A$ over $B$ (Pareto). Arrow proved that if there are at least three alternatives, then the only aggregation functions satisfying these desiderata are dictatorships: there is one individual $d$ such that no matter what others prefer, if $d$ strictly prefers $A$ to $B$, then the social preference ranks $A$ over $B$ as well. [^2] Mishra (2023) applies Arrow’s Theorem to RLHF. A similar theorem (see Taylor 2005, § 1.3) holds for social choice functions where, instead of asking for a social ranking of alternatives, we more modestly ask for just a set of choice-worthy alternatives. This also includes the special case of social choice functions that always pick a single winner.
Arrow’s Theorem stimulated a huge literature exploring the consequences of weakening Arrow’s desiderata (see, e.g., Campbell & Kelly 2002, Holliday & Pacuit 2020, and references therein). The general takeaway is that for ordinal preference aggregation, in order to avoid dictatorships and related pathologies such as oligarchies and vetoers, one must weaken the Independence of Irrelevant Alternatives (IIA) and allow the social preference between two alternatives to depend in part on individual preferences involving other alternatives. With this freedom to relax IIA comes a vast proliferation of alternative methods of aggregating individual preferences (see, e.g., Brams & Fishburn 2002; Zwicker 2016; Pacuit 2019 and the voting methods implemented in the Preferential Voting Tools library). Figure 1 gives an example in which three well-known methods disagree. The costs and benefits of these and other methods are systematically studied from different angles (axiomatic, computational, empirical, etc.) in social choice theory.
$4$ $4$ $9$ $4$ $2$
$A$ $A$ $B$ $C$ $C$
$B$ $C$ $C$ $A$ $B$
$C$ $B$ $A$ $B$ $A$
Borda Count: $CBA$
Instant Runoff: $ABC$
Ranked Pairs: $BCA$
Figure 1: Individual rankings on the left (4 voters submit the ranking $ABC$, 4 submit $ACB$, etc.) lead to different aggregated rankings on the right, depending on the aggregation rule. Borda Count gives an alternative 0 points for each voter who ranks it last, 1 point for each voter who ranks it second, and 2 points for each voter who ranks it first; alternatives are then ordered by descending score. Instant Runoff ranks $C$ last since $C$ has the fewest first-place rankings; then, after removing $C$ from all voters’ rankings, $B$ has the fewest first-place rankings, so $B$ is in second and $A$ is in first. For Ranked Pairs, notice there is a majority cycle: a majority of voters prefer $A$ to $B$, a majority prefer $B$ to $C$, and a majority prefer $C$ to $A$; but the smallest majority margin of victory is for $A$ over $B$, so we reverse this majority preference, yielding $BCA$.
Since Arrow, social choice theory has grown to study aggregation not only of individuals’ preferences, both ordinal and cardinal (d’Aspremont & Gevers, 2002), but also of their approvals of alternatives (Laslier & Sanver, 2010), grades given to alternatives (Balinski & Laraki, 2010), judgments about propositions (Grossi & Pigozzi, 2022), subjective probabilities for propositions (Dietrich & List, 2016), and other types of objects (Rubinstein & Fishburn, 1986). In the following, we discuss some of the aggregation problems that might arise in the context of AI alignment.
3 What are the Collective Decision Problems and their Alternatives in this Context?
If we want to use methods from social choice for the purpose of aligning AI systems, we first need to specify what the concrete options/objects are, before we can start collecting preferences over them and make actual or simulated collective choices between them. These options are called alternatives in social choice theory. In some contexts, the set of alternatives is easy to comprehend and enumerate, as when the alternatives are candidates for a position or an award. In other settings, there are exponentially many alternatives, but the set is still easy to comprehend, e.g., when there are $n$ propositions and each of them must be either accepted or rejected (Lang, 2007).
When considering the alignment of AI systems, it is harder to see exactly how best to think about the relevant set of alternatives for evaluation. In principle, it could be the set of all AI systems or all possible parameterizations of a given network architecture, but this would surely be conceptually intractable.
In the context of an LLM, the RLHF approach traditionally asks the evaluator to choose between a small, explicit set of alternative responses to a single prompt, with each response sampled from the LLM’s output distribution. Alternately, we could consider all possible responses as alternatives. While this response set is too large to explicitly enumerate, the evaluators can still indicate their preference by providing the preferred response themselves. Such exemplars are often used for fine-tuning and can be used to learn evaluators’ preferences and generate responses that well-represent them (Fish et al., 2023). While this does not address questions about how to generalize beyond a single prompt, it is a useful way of conceptualizing the alternatives.
One might conceive of the alternatives as probability distributions over responses. This is natural, as LLMs are typically configured to respond stochastically to a prompt. This might be desirable not only for creativity but also to promote fairness and representativeness of responses. For example, in response to a controversial question, fairness might militate against an LLM always giving the same answer, as any one answer will inevitably omit some relevant considerations on one side of a debate. There is a large literature on social choice rules whose outputs are probability distributions. The inputs to such a rule could be the evaluators’ stated explicit preferences between distributions (Fishburn 1973, Ch. 18), but they could also be stated preferences between plain alternatives (Brandt, 2017). Indeed, the type of objects chosen by a social choice rule (e.g., distributions over responses) need not match the type of objects about which individuals state their preferences or evaluations (e.g., responses). This is important, since probability distributions over large sets of responses may be particularly difficult for evaluators to reliably compare.
A multi-winner rule (Faliszewski et al., 2017; Elkind et al., 2017) could be a middle ground between a deterministic single-winner social choice rule and a probabilistic rule. Such a rule picks a small, predetermined number of answers that best reflect what the voters want. These winning answers could then be combined into a single response that lists these winning answers as bullet points to provide the user with a representative overview of possible answers.
4 Who Provides the Human Feedback?
Let us assume that there is a population of people, the stakeholder population, who will be affected by an AI system and whose preferences would therefore ideally be taken into account in aligning the AI system. [^3] There may also be stakeholders, such as small children and non-human animals, whose feedback we cannot easily elicit. In that case, we may consider feedback from humans who are charged with representing their interests. Unfortunately, it may be infeasible to elicit feedback from all members of the stakeholder population, so we must select some smaller group from which to elicit feedback. For example, one could try to select a suitably representative subset of the population such that the alignment obtained using feedback from the subset sufficiently approximates the alignment that would be obtained using feedback from the full stakeholder population. Here one could draw on ongoing work in social choice theory on how to select citizens’ assemblies that are representative of a full population (e.g., Flanigan et al. 2021; Landemore & Fourniau 2022), as well as work in statistics on efficient stratified sampling (e.g., Meng 2013).
Another approach would be to allow the full stakeholder population to vote on their representatives in some way. This could be done, for example, with a voting procedure that is designed to elect assemblies that are proportionally representative (see, e.g., Ch. 4 of Lackner & Skowron 2023). Additionally, stakeholders might be allowed to delegate their feedback rights to others (who may in turn delegate, etc.), as in liquid democracy (see Paulin 2020).
As of now, earlier work has used evaluator recruitment methods such as Mechanical Turk (Freedman et al., 2020; Bai et al., 2022a); Upwork, Scale AI, or Lionbridge (Stiennon et al., 2020; Ziegler et al., 2019); and purpose-built platforms (Noothigattu et al., 2018). We believe this component of the RLHF pipeline deserves a more in-depth discussion, including one informed by social choice theory.
5 What is the Format of Human Feedback?
As we have discussed, human feedback for AI systems can come in various forms; which of these are most natural and useful? Here, we can draw on a significant literature on preference elicitation (see, e.g., Sandholm & Boutilier 2006), studying how best to query agents for their preferences in a variety of domains. [^4] Incidentally, recently, it has been proposed to use LLMs as a tool to help do so (Li et al., 2023). This literature is closely tied to that of communication complexity (e.g., Kushilevitz & Nisan 1997), which is concerned with minimizing the number of bits that need to be communicated to achieve something. (Though the number of bits communicated is of course generally not the perfect way to measure how much effort a human participant needs to expend to answer queries.) Preference elicitation and communication complexity have also been studied in voting settings (Conitzer & Sandholm, 2002, 2005; Service & Adams, 2012).
5.1 Multiple Format Options
In general, we want the type of input or feedback that we ask of humans to be (1) natural to give, (2) informative about their preferences and values, and (3) of a type that can be used to align AI systems. For example, with current methods, having humans comment on an AI output in an open-ended text box may satisfy 1 and 2, but not 3. Having them sort responses alphabetically may satisfy 1 and 3, but not 2. Having them directly rank neural networks based on inspecting their weights may satisfy 3 but not 1 or 2.
It should be noted that different choices for the type of input or feedback can lead to differently aligned systems, especially if we do not understand the behavioral effects of the different types of input. For example, McElfresh et al. (2021) introduce (in the context of feedback on kidney allocation) an indecision option among the available choices and reject several natural hypotheses about how the resulting data relate to those obtained without that option.
One question is whether we should actually let individual humans choose the format in which they give input or feedback. In traditional social choice, this is uncommon, although there may be some flexibility in how preferences are expressed or input is given (e.g., allowing voters to not give a complete ranking but rather only rank a few alternatives (Halpern et al., 2023), or to give numerical ratings instead of ordinal rankings), as well as some variety in the interaction mechanism to get to that expression of preferences (e.g., one can vote for candidates individually but also pull a lever that corresponds to voting for exactly the candidates of a single party).
It is easy to imagine giving evaluators the choice between a range of different ways to give their input or feedback on various aspects of the system’s behavior or behavioral patterns or rules (e.g., individual responses, whole dialogue sessions, longterm interaction with the same user, or published guiding principles) and various dimensions of desirability, which is emerging as fine-grained RLHF (Wu et al., 2023) or optimizing attributes in the data (Dong et al., 2023), relating to various values such as “truthfulness”, “harmlessness”, “fairness”, etc., and to allow them to give that feedback in various ways: approving/disapproving, making pairwise comparison statements of the form “I like A better than B”, giving full or partial rankings of the form “A is best, B 2nd-best, …”, giving precise or imprecise ratings of the form “I rate A between 7 and 9”, or even by giving free-form verbal feedback that the LLM then interprets and converts into some formal data such as a partial ordering. This heterogeneous data could then be transformed in some formal way into a common, sufficiently expressive data structure, such as a utility function.
5.2 Dealing with Diverse and Informal Feedback
Recall that in RLHF, human feedback is typically used to train a reward (or “preference”) model whose job it is to map any possible AI system response to a numerical rating. The concept of reward models could also be used to convert the diverse input or feedback of a single evaluator into a common form, in order to then aggregate it with the input of other evaluators to steer an AI system.
First, an individual evaluation interpretation model $\phi$ could be trained to map a tuple of inputs of the form $(x,\mathcal{Y},f_{i},e,y)$ to a numerical evaluation $r$. As before, $x$ represents a prompt to the AI system, $\mathcal{Y}$ the set of possible AI responses, and $y\in\mathcal{Y}$ a particular response. Moreover, vector $f_{i}$ represents the relevant features of a certain evaluator $i$, and $e$ shall be a language representation of $i$’s feedback on possible responses $\mathcal{Y}$ to $x$, containing preference- and evaluation-related statements of whatever type (see Section 5.1). In practice, $\phi$ would likely be based on an LLM pretrained to understand the texts $x$, $\mathcal{Y}$, $e$, and $y$, that is then fine-tuned to the interpretation task described above. Then the output $r=\phi(x,\mathcal{Y},f_{i},e,y)$ of $\phi$ is a numerical rating of $y$ given by evaluator $i$ that is trained to be (approximately) consistent with the verbal evaluation $e$ of that evaluator. We note that this task can be seen as a form of meta-learning.
One could then use the trained evaluation interpretation model $\phi$ to train another model—an individual preference model $\psi$—that skips verbal evaluations and directly maps inputs $(x,\mathcal{Y},f_{i},y)$ to ratings $r=\psi(x,\mathcal{Y},f_{i},y)$. Namely, any tuple $(x,\mathcal{Y},f_{i},e)$ can be converted into supervised training data $\big{(}(x,\mathcal{Y},f_{i},y),\phi(x,\mathcal{Y},f_{i},e,y)\big{)}_{y\in\mathcal{Y}}$ for $\psi$, containing simulated ratings $r=\phi(x,\mathcal{Y},f_{i},e,y)$. The hope is that the individual preference model $\psi$ would be able to simulate the rating of any evaluator (represented by their features $f_{i}$), as long as the evaluator, prompt, and response set come from the same distribution as the one $\psi$ was trained on. Similar to the preference models used in current RLHF, $\psi$ could finally be used to fine-tune the actual AI system or steer its behavior in real time. In fact, if the evaluators’ features $f_{i}$ were omitted in the training process sketched above, $\psi$ would be a preference model of the same type as is already used in RLHF and could readily be used for it. This would, however, conflate the evaluations of the (possibly not proportionally representative) set of evaluators used in training in a rather uncontrolled and potentially confusing way. An arguably better way of making use of $\psi$ is, therefore, to indeed make use of evaluators’ features $f_{i}$ in training and add an additional social choice step to the RLHF pipeline or the AI system’s real-time decision-making procedure. Below we sketch several ways in which this might be done.
6 How can Diverse Individual Input or Feedback be Incorporated?
Here we sketch several variants of two approaches for including diverse input or feedback into AI systems in a consistent way using methods from social choice theory. The first suggests adding an additional preference aggregation step somewhere during training, thereby turning RLHF into RLCHF: Reinforcement Learning from Collective Human Feedback. The second approach instead suggests adding an additional simulated collective decision step somewhere in the training or the system’s real-time decision procedure, similar to Bakker et al. (2022) and Jarrett et al. (2023).
Figure 2: RLCHF using aggregated rankings. The core addition to the standard RLHF process is the call-out of an explicit social welfare function, $F$, which determines how preferences are aggregated.
6.1 Proposal: Reinforcement Learning from Collective Human Feedback (RLCHF)
Preference aggregation could be incorporated as an additional step into RLHF in several ways, from early to rather late in the RLHF pipeline. For clarity of exposition, assume a simple version of rankings-based RLHF that (1) takes a database of prompts $x$ together with corresponding sets of possible responses $\mathcal{Y}$, (2) asks one associated evaluator $i(x,\mathcal{Y})$ to provide a ranking $R(x,\mathcal{Y})$ of the elements of $\mathcal{Y}$, (3) turns this ranking into $|\mathcal{Y}|$ many data points for training a common preference model $\varrho$ that produces numerical ratings $r=\varrho(x,y)$, and (4) uses these ratings as rewards in fine-tuning the actual LLM via reinforcement learning.
The earliest point to introduce preference aggregation in this pipeline would be between steps (2) and (3). Instead of a single evaluator $i(x,\mathcal{Y})$, we may ask the members of a jury $J(x,\mathcal{Y})$ of evaluators to provide individual rankings $R_{j}$. Using some ordinal social welfare function $F$, those rankings can then be aggregated into a collective ranking $R=F((R_{j})_{j\in J})$ to use it in step (3). This approach could be termed “RLCHF using aggregated rankings” and is shown in Fig. 2.
Figure 3: RLCHF using evaluator features and aggregated ranks. We show how an individuals’ features can be used as an additional input to reward models within the RLHF process.
Alternatively, one could use cardinal rather than ordinal preference aggregation at a later point in the pipeline: between steps (3) and (4). For this, change step (3) so that not a model of common but of individual preferences is trained, mapping pair $(x,\mathcal{Y})$ and evaluator $i$ with features $f_{i}$ to predicted ratings $r_{i}=\psi(x,f_{i},y)$. Also generate a large collection of feature vectors $f_{1},\dots,f_{N}$ that is representative of the stakeholder population. Then a cardinal social welfare function $W$ can be used to aggregate into one rating $\varrho(x,y)=W(\psi(x,f_{1},y),\dots,\psi(x,f_{N},y))$ which can be used in step (4). This approach could be termed “RLCHF using evaluator features and aggregated ratings” and is shown in Fig. 3.
6.2 Proposal: Simulated Collective Decisions
RLCHF, as described above, keeps the reinforcement learning step that requires numerical rewards, and it uses ordinal or cardinal preference aggregation to produce these said rewards for all possible responses $y\in\mathcal{Y}$. A different approach would replace reinforcement learning by something else and introduce social choice methods in the form of simulated collective decisions rather than preference aggregation.
For one thing, one could modify “RLCHF using evaluator features and aggregated ratings” into “Supervised Learning from Simulated Collective Decisions”, as shown in Fig. 4. For this, in step (3) from above, use the individual preference model $r_{i}=\psi(x,f_{i},y)$ and feature vectors $f_{1},\dots,f_{N}$ not to produce an aggregated rating but to simulate a collective choice that picks a single winning response $y^{*}=C\big{(}(\psi(x,f_{j},y))_{y\in\mathcal{Y},j=1,\dots,N}\big{)}$. Here, $C$ is now a single-winner social choice function. Then in step (4), use data point $(x,y^{*})$ to train the actual AI system via supervised (rather than reinforcement) learning. Instead of picking a single winner $y^{*}$, we could also use a multi-winner social choice function $C$ that outputs, say, a set of three responses $(y^{\prime},y^{\prime\prime},y^{\prime\prime\prime})$. These can then be (creatively) combined into a single response, for example, by merging them into a bullet-point list and adding a sentence “The following are (three) typical answers to your question: …” at the beginning.
A more radical modification would drop the fine-tuning-via-learning step altogether (leaving the LLM only pretrained) and rather simulate the collective choice at inference time. Whenever the live system is prompted with some $x$, generate $k\gg 1$ many candidate responses $y_{i}$ and $N\gg 1$ many evaluator feature vectors $f_{j}$ representative of the stakeholder population for the problem $(x,\mathcal{Y})$, and directly return the winner $y^{*}=C\big{(}(\psi(x,f_{j},y_{i}))_{j,i=1}^{N,k}\big{)}$ of the simulated collective choice. Here, too, $C$ could be a multi-winner or probabilistic social choice rule.
7 Which Traditional Social-choice-theoretic Concepts are Most Relevant?
A wide variety of concepts is studied in social choice. In general, the relevance of those concepts depends on the specific application. For example, consider the concept of false-name-proofness (Yokoo et al., 2001, 2004; Conitzer & Yokoo, 2010), which means that no participant can benefit from participating multiple times under multiple accounts. This concept is relevant when voting over the internet, but entirely irrelevant in, say, an in-person faculty meeting where faculty vote publicly by raising their hands.
So, rather than studying every single social-choice-theoretic concept in the context of aligning AI systems, we should be careful to evaluate which traditional concepts are most relevant. In the following, we give just a few examples.
7.1 Independence of Clones
Figure 4: Supervised Learning from Simulated Collective Decisions. We show that with an individual or cardinal reward model, as presented in Figure 3, responses $y$ to a prompt $x$ can be simulated. This process expands the scope of studying preferences within RLHF and opens future work on personalization and other topics.
In social choice problems, sometimes multiple alternatives, say $A$ and $B$, compare very similarly against every other alternative $X$, according to the preferences of individuals. Such alternatives are referred to as clones, a notion that can be formalized in several ways. According to a strict notion of clones (Tideman, 1987), $A$ and $B$ are clones if, for every individual, if that individual prefers $A$ to some other alternative $X$, then they also prefer $B$ to $X$, and if they instead prefer $X$ to $A$, then they also prefer $X$ to $B$. According to a more liberal notion (Laffond et al., 1996), $A$ and $B$ are clones if, whenever a majority of individuals prefer $A$ to some other alternative $X$, then a majority prefers $B$ to $X$ as well, and whenever a majority prefers some $X$ to $A$, then a majority prefers $X$ to $B$ as well.
Sometimes the introduction of a clone can affect the outcome of an election. For example, suppose a group of people are voting over where to go for dinner, and the only two alternatives are a Chinese restaurant and an Indian restaurant. 52% of the voters prefer the Chinese restaurant. But then, someone points out that the Chinese restaurant has two floors and argues that the two floors should be considered separate options. So now the alternatives are $C_{1}$, $C_{2}$, and $I$. It turns out nobody really cares all that much about the floor, but suppose that 26% of the voters prefer $C_{1}\succ C_{2}\succ I$, and 26% of the voters prefer $C_{2}\succ C_{1}\succ I$ (adding up to the original 52%). Further suppose that the voting rule used is Plurality, in which the alternative that appears at the very top of voters’ rankings the most often wins. [^5] Given that only the top-ranked alternative matters, in practical implementations of this rule, we simply ask voters to list only that alternative—the alternative the voter votes for. This results in the Indian restaurant now actually winning with 48% of the vote. This seems like an undesirable property for a voting rule to have; it would be better for the introduction of a clone never to make a difference. [^6] More precisely, introducing a clone should not affect whether a non-clone (e.g., the Indian restaurant in our example) is selected or which non-clone is selected. But it may affect which clone, if any, is selected (e.g., a clone-independent rule could select $C_{1}$ over $C_{2}$ in our example, if among the 48% of people who prefer $I$, a strict majority of them prefer $C_{1}$ to $C_{2}$). This latter desirable property is called independence of clones. Perhaps when choosing restaurants, this is not that important, as restaurants will rarely be clones (unless the floors of restaurants are treated separately). On the other hand, when choosing responses for a chatbot, it may be quite common for two responses to be very close to each other. This suggests that this property is more important in this context.
7.2 Strategic voting
Another concern is strategic voting (or strategic feedback). Strategic voting consists of casting a vote that does not reflect one’s true preferences, in order to obtain a better result for oneself. For example, consider an election with plurality voting, as described above. A voter might perceive that her top-ranked alternative has no chance of winning and therefore strategically vote for another alternative. Strategic voting poses a problem because we can no longer take votes (or feedback) at face value. Unfortunately, in general, every reasonable voting rule will sometimes introduce incentives to manipulate (Gibbard, 1973; Satterthwaite, 1975). These incentives to manipulate might be reduced if voters lack full information about the preferences of other voters (Conitzer et al., 2011) or about the voting rule that will be used (Holliday & Pacuit, 2019). But we often cannot guarantee such ignorance, just as we often cannot guarantee computer security through obscurity.
What form might strategic voting in a context such as RLHF take? If rating responses on a scale from (say) 0 to 10, a natural strategy is to overreport. E.g., if one evaluator does not really like a response (at the level of a 3), but suspects that others would like it (say, two other evaluators that give a 6), then this evaluator may strategically give a rating of 0 to “compensate” for the other reviewers. This manipulation would be successful if we eventually aggregate ratings by taking their average: the average will be pulled down to 4, instead of the 5 that would result from reporting truthfully, so that the average is closer to the 3 that the evaluator believes is ideal. If instead we use the median as the aggregate, then this manipulation is ineffective—the median would remain 7. Indeed, the median is strategy-proof in this context: misreporting one’s preferences never helps, as long as one’s only goal is to move the median rating closer to one’s “true” rating.
7.3 Anonymity
In democratic contexts, a standard desideratum on voting rules is anonymity: if two voters swap their ballots before submitting them, the output of the voting rule will not change (the rules in Figure 1 all satisfy anonymity in this sense). This captures the idea that the voting rule should not favor some voters over others. Anonymity not only prohibits the extremes of dictatorship (recall Section 2.5), but even any kind of weighted voting wherein some voters’ votes count for more than others. However, in the context of AI development, one might consider aggregating human feedback in a way that violates anonymity (cf. the weighted majority rule discussed in Nitzan & Paroush 1982). Perhaps some evaluators are more experienced or more highly rated than others; perhaps some are influenced by others, so their input should not be considered completely independent inputs for aggregation; and so on. In general, whether the same democratic norms applied to voting also apply in an AI context is an important question for discussion.
7.4 Principles as Voters
While it is standard in social choice for the voters to be human agents, this is not the only interpretation of the mathematical framework of social choice theory. In some applications of social choice to AI ethics and safety, possibly including Constitutional AI (recall Section 2.4), we might regard different ethical principles as the “voters” who can rank or otherwise evaluate the outputs of an AI system. (cf. Greene et al. 2016.) This is analogous to applications of social choice theory in the philosophy of science, where the “voters” are theoretical virtues that may rank scientific theories differently (Okasha, 2011), or to multi-criteria decision-making, where the “voters” are relevant factors that may rank the options differently (Arrow & Raynaud, 1986). Of course, such ethical principles could themselves be outputs of some prior social choice procedure in which the voters are humans (cf. Collective Constitutional AI in Ganguli et al. 2023).
This principles as voters idea suggests a possible alternative architecture for applying social choice to AI—one sitting somewhere between the extremes of a spectrum that ranges from Constitutional AI at one end (in which principles are the whole show, while social choice does not appear) to the RLHF version of reinforcement learning as described above (in which principles play no role at all). In this alternative model, each respondent would be required to justify her rankings of alternative AI responses in terms of their level of satisfaction of each of a number of principles taken from a fixed menu. The AI system would use the results to train for several independent tasks: for each principle, separately learn how to rate responses to queries based on that principle alone; and learn how to aggregate those separate ratings into an overall rating of the responses. These would be composed to form the final stage of a simulated collective decision—the stage in which the voters are the principles.
8 How Should We Account for Behavioral Aspects and Human Cognitive Structures?
The theory of preference elicitation tends to be based on idealized assumptions; for example, each agent whose preferences are being elicited has well-defined and consistent preferences, and the agent answers in a way that is perfectly consistent with those preferences (or perhaps random noise is added). In reality, when we elicit human preferences, myriad behavioral effects kick in.
This leads to a variety of questions about how to align AI systems in the context of such behavioral effects. Should we correct for the behavioral effect? That would seem to require having a model where such a behavioral effect obscures humans’ “true” values. But do these “true” values correspond to anything real in the world? Do we run the risk of the “correction” actually removing valuable information? Could the ability to make such “corrections” in fact be abused to intentionally remove inconvenient feedback?
9 How to Navigate a Multiplicity of AIs?
Consider the example of a group of people voting over the restaurant where they will go for dinner. If there is significant disagreement in the votes, rather than forcing a minority to go to a restaurant that they really do not like, it can make sense to split the people into multiple groups, each going to their own restaurant. Similarly, in the current context, perhaps it makes sense to create multiple AI systems; for example, to recognize strong inter- and intra-cultural variations that have been identified in some non-homogenous populations (Awad et al., 2018; Peters & Carman, 2024). Depending on the situation, the people providing feedback might be split into groups ex ante (for example, country A makes one system based on the feedback of A’s citizens and country B another based on the feedback of B’s citizens), but also ex post, where we first collect feedback and then consider which people it makes sense to group together. The latter approach is closely related to the topic of representation in voting theory (Faliszewski et al., 2017).
There is also the slightly different scenario where one AI system is in place, and some group of people believes that it is not serving them well. Hence, they might decide to pool their resources and create their own system. The literature on cooperative game theory (cf. Chalkiadakis et al. 2011), sometimes referred to as coalitional game theory, touches on these considerations (and indeed also plays a role in questions of representation, Aziz et al. 2017).
Finally, let us highlight possible shortcomings to creating multiple AI systems. As in the restaurant example, it may have the result of unnecessarily dividing people into separate groups. Moreover, splitting into groups may not be feasible if it does not dovetail with existing social structures. For example, the US Federal Government may want to adopt a single system that will impact all its citizens, and adopting two systems would be tantamount to splitting the country in two. Finally, unlike in the case of the restaurants, the multiple AI systems may have to interact with each other, creating the risk of conflict between AIs with different goals. The nascent literature on cooperative AI (Dafoe et al., 2021; Conitzer & Oesterheld, 2023) may help keep these kinds of interactions from going horribly wrong. Nonetheless, it might be best to see if we can completely avoid having multiple AIs with competing goals, or at least design them in a way that makes conflict between them less likely.
10 Conclusion
It is important that a variety of stakeholders are involved in giving input or feedback on how AI systems, such as those based on LLMs and other foundation models, should function. But those stakeholders are likely to give conflicting input. If so, how do we aggregate this input or otherwise use it for real or simulated collective decisions to end up with a sensible system? As we have argued in this paper, the field of social choice is well placed to help address this question—conceptually, due to its focus on methods for making consistent collective decisions, e.g., via aggregating preferences, judgments, and other inputs in a consistent way, as well as pragmatically, with many researchers in the computational social choice community being well prepared to engage with AI alignment researchers on these problems.
That said, it is important to acknowledge that aggregating conflicting input or feedback can be a complex task. It requires careful consideration of various factors, such as who the stakeholders are, which humans should provide the feedback, how their input is collected and weighed, the level of expertise and credibility of their input, and potential biases. Additionally, incorporating transparency and accountability measures into the aggregation process can help ensure that the final system reflects a fair and balanced representation of the stakeholders and their input. Significant research is needed to deepen our understanding of the possibilities of using social choice for these purposes and the different effects that this will have.
Needless to say, the questions considered above are multifaceted and, as such, cannot be adequately addressed without complementary (not necessarily AI-specific) research. How best to make practical decisions, as well as associated legal and political considerations, provide further important avenues for future research.
Last but not least, we have put a particular focus on RLHF in this paper as it is an especially important and fruitful point of contact between social choice and AI. But the insights afforded by social choice theory bear on countless problems. Social choice can be used to more generally determine the objectives that AI systems pursue, the data on which they are trained, and which systems we build in the first place. Given the rapid development of AI systems underway, we urge researchers to begin forging these connections between social choice and AI alignment.
Acknowledgements
This paper grew out of the workshop on Social Choice for AI Ethics and Safety held at UC Berkeley in December 2023. We thank all the participants of the workshop for fruitful discussions. We are also grateful to Open Philanthropy for a grant that made the workshop possible, as well as further support from the Center for Human-Compatible AI, the C3.ai Digital Transformations Institute, and the Kavli Center for Ethics, Science, and the Public. For helpful comments on this paper, we thank Aditi Raghunathan and anonymous reviewers.
Impact Statement
This paper highlights the need for further research and collaboration between experts in social choice and AI ethics and safety to ensure that AI systems are designed and deployed in a way that aligns with societal values and promotes accountability and transparency. As we discussed briefly in the introduction, we believe the approach proposed in this paper would result in systems that are fairer, that are less likely to have blind spots due to nobody having raised an issue, and that people buy into more broadly.
As we briefly discussed in the introduction, there is perhaps a concern that feedback from a broader set of participants is more likely to be inconsistent and that consequently the resulting system will behave erratically. The idea that naïve aggregation of votes or judgments leads to inconsistency is a familiar one from social choice theory. For example, if three voters rank three alternatives $A$, $B$, and $C$ respectively as follows: $A\succ_{1}B\succ_{1}C$, $C\succ_{2}A\succ_{2}B$, and $B\succ_{3}C\succ_{3}A$, then a majority of voters prefers $A$ to $B$, a majority prefers $B$ to $C$, and a majority prefers $C$ to $A$. This illustrates that majority preferences are cyclical and thus arguably irrational. (See Figure 1 for another example.) We encounter similar issues in judgment aggregation (for an overview, see Endriss 2015). To illustrate this in our own context, say that it is broadly agreed that an output should be given if and only if it is both safe and helpful. Suppose evaluator 1 believes the output is safe but not helpful, and therefore should not be given. Evaluator 2 believes the output is helpful but not safe, and therefore should not be given. Evaluator 3 believes the output is both safe and helpful, and therefore should be given. Then a majority believes that the output is safe, a majority believes that it is helpful, but a majority believes that it should not be given—so that majority judgments are logically inconsistent. However, social choice theory is precisely concerned with how we should actually obtain consistent aggregations and therefore is well placed to address this issue. For example, one common strategy is to restrict to rational or consistent outputs only and among these find one that is in some sense “closest” to the reports (see, e.g., Elkind & Slinko 2015). Therefore, social choice theory is actually well positioned to help with the issue of inconsistencies from aggregation.